OpenStudy (anonymous):

Simplify (x^k + 1)^2(–x^k)^3(3x).

OpenStudy (anonymous):

PLEASE help anyone!

OpenStudy (anonymous):

ok well first what can you do with the first part, \[(X ^{k}+1)^{2}\]

OpenStudy (anonymous):

how do you square binomials

OpenStudy (anonymous):

Using the binomial formula... right?

OpenStudy (anonymous):

so what does that look like?

OpenStudy (anonymous):

is this what the whole thing looks like\[(x ^{k}+1)^{2}(-x ^{k})^{3}(3x)\]

OpenStudy (anonymous):

No it looks differently. It looks like this (x ^k + 1) ^2(–x^k) ^3 (3x)

OpenStudy (anonymous):

ok so what do you do with the first part (x^k+1)^2

OpenStudy (anonymous):

I am not sure but, is it x^k+2 ?

OpenStudy (anonymous):

not exactly, this is a hard problem, when you square binomials you use foil, have you learned that yet?

OpenStudy (anonymous):

Yes I have, but I am not sure how I would use it in this equation :(

OpenStudy (anonymous):

well you know how 2^2 = 2 times 2

OpenStudy (anonymous):

No this sign ^ means it is in the power of... not times

OpenStudy (anonymous):

I am sorry for confusing you

OpenStudy (anonymous):

I understand, 2 was not a good example. how about 8^2 isn't that 8 times 8

OpenStudy (anonymous):

Yes, I understand now

OpenStudy (anonymous):

well when you square a binomial you do the same thing, so what do you think it will look like?

OpenStudy (anonymous):

x^k^2+2

OpenStudy (anonymous):

not quite, that is the most commonly missed problem. write it out (x^k+1)^2= (x^k+1)(x^k+1)

OpenStudy (anonymous):

Oh yes now I get it. Thanks

OpenStudy (anonymous):

so What do you do next

OpenStudy (anonymous):

you simplify the other (-x^k)^3

OpenStudy (anonymous):

well first foil (x^k+1)(x^k+1) this is tricky

OpenStudy (anonymous):

I don't know how to continue

OpenStudy (anonymous):

do you know what it means to FOIL?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is it x^2^k^2+1

OpenStudy (anonymous):

let's slow down, well what is the first step, what do you do with the first terms

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

x^k is the first term so how do you multiply the first terms